# Stochastic gradient descent estimation of generalized matrix factorization models with application to single-cell RNA sequencing data

### v1

@@ 3.2.1 Improving naïve stochastic gradient
where $\bar{\mathbf{G}}_{\psi}^{t}$ and $\bar{\mathbf{H}}_{\psi}^{t}$ are smoothed estimates of the derivatives of (6), while $\alpha_{t}$ is a scalar bias-correction factor. Similarly to Adam (Kingma and Ba,, 2014) and AMSGrad (Reddi et al.,, 2019), we update the smoothed gradients, $\bar{\mathbf{G}}_{\psi}^{t}$ and $\bar{\mathbf{H}}_{\psi}^{t}$, using an exponential moving average of the previous gradient values and the current stochastic estimates $\hat{\mathbf{G}}_{\psi}^{t}$ and $\hat{\mathbf{H}}_{\psi}^{t}$, namely
@@
with $\alpha_{1},\alpha_{2}\in(0,1]$ being user-specified smoothing coefficients; $\alpha_{t}=(1-\alpha_{2}^{t})/(1-\alpha_{1}^{t})$ is introduced to filter out the bias induced by the exponential moving average in (12). Similar smoothing techniques have two main advantages: they speed up the convergence using the inertia accumulated by previous gradients and, at the same time, they stabilize the optimization reducing the noise around the gradient estimate. We refer the reader to Kingma and Ba, (2014) and Reddi et al., (2019) for a deeper discussion on the benefits of bias corrected exponential gradient averaging in high-dimensional stochastic optimization problems.
Differently from the original implementation of Adam (Kingma and Ba,, 2014), which only involves first-order information, we scale the smoothed gradient estimate using a smoothed diagonal Hessian approximation. This strategy allows us to directly extend the diagonal quasi-Newton algorithm of Kidziński et al., (2022) in a stochastic vein without increasing the computational overload or the stability of the optimization. Indeed, the derivatives of the deviance function of a GLM model are always well-defined, bounded away from zero, available in closed form and easy to compute. This is not the case for deep neural networks and other machine learning models for which Adam and AMSGrad have been originally developed.
Similarly to representation (10) for GMF models, also (11) can be written in the block-wise form
@@
In this formulation, update (10) and (13) have a computational complexity proportional to the dimension of the matrices $\mathbf{U}$ and $\mathbf{V}$, namely $O((n+m)d)$. Since the dimension of the parameter space can not be further compressed without losing prediction power, the alternative way to speed up the computations is to find an efficient way for approximating the gradients, $\hat{\mathbf{G}}_{\psi}^{t}=(\hat{\mathbf{G}}_{\textsc{u}}^{t},\hat{\mathbf{G}}_{\textsc{v}}^{t})$ and $\hat{\mathbf{H}}_{\psi}^{t}=(\hat{\mathbf{H}}_{\textsc{u}}^{t},\hat{\mathbf{H}}_{\textsc{v}}^{t})$.
@@
In what follows, we discuss a new stochastic optimization method for GMF models based on formula (13), which use local parameter updates along with minibatch estimates of the current gradients to reduce the computational complexity of the resulting algorithm. To this end, we introduce the following notation: $I=\{i_{1},\dots,i_{n_{{\scriptscriptstyle{I}}}^{\star}}\}\subseteq\{1,\dots,n\}$ is a subset of row-indices of dimension $n_{{\scriptscriptstyle{I}}}^{\star}$, $J=\{j_{1},\dots,j_{m_{{\scriptscriptstyle{J}}}^{\star}}\}\subseteq\{1,\dots,m\}$ is a subset of column-indices of dimension $m_{{\scriptscriptstyle{J}}}^{\star}$, $B=I\times J$ is the Cartesian product between $I$ and $J$. Finally, $\mathbf{Y}_{{\scriptscriptstyle{I}}:}=\{\mathbf{y}_{i:}\}_{i\in I}$, $\mathbf{Y}_{:{\scriptscriptstyle{J}}}=\{\mathbf{y}_{:j}\}_{j\in J}$ and $\mathbf{Y}_{{\scriptscriptstyle{B}}}=\{y_{ij}\}_{(i,j)\in B}$ denote the corresponding sub-matrices, also called row-, column- and block-minibatch subsamples of the original data matrix, respectively.
@@ 3.2.2 Block-wise adaptive stochastic gradient descent
Both the exact and stochastic quasi-Newton methods identified by equations (10) and (13) entirely update $\mathbf{U}$ and $\mathbf{V}$ at each iteration. Despite being an effective and parallelizable strategy, in many applicative contexts, when it comes to factorizing huge matrices, an entire update of the parameters could be extremely expensive in terms of memory allocation and execution time. This is a well-understood issue in the literature on recommendation systems, where standard matrix completion problems may involve matrices with millions of rows and columns; see, e.g., Koren et al., (2009), Mairal et al., (2010), Recht and Ré, (2013), Mensch et al., (2017). Moreover, batch optimization strategies do not generalize well to stream data contexts, where the data arrive sequentially and the parameters have to be updated on-the-fly as a new set of observations comes in.
A classic solution is to perform iterative element-wise SGD steps using only one entry of the data matrix, $y_{ij}$, at each iteration, thus updating the low-rank decomposition matrices row-by-row through the paired equations $\mathbf{u}_{i:}^{t+1}\leftarrow\mathbf{u}_{i:}^{t}-\rho_{t}\mathbf{G}_{\textsc{u},i:}^{t}$ and $\mathbf{v}_{j:}^{t+1}\leftarrow\mathbf{v}_{j:}^{t}-\rho_{t}\mathbf{G}_{\textsc{v},j:}^{t}$. In the matrix factorization literature, this strategy is also known as online SGD algorithm, since it permits to update the parameter estimates dynamically when a new observation is gathered.
To stabilize the optimization and speed up the convergence, we generalize the online SGD approach in two directions: we consider local stochastic quasi-Newton updates in place of naïve gradient steps and we use block-wise minibatches of the original data matrix, $\mathbf{Y}_{\scriptscriptstyle{B}}$, instead of singletons, $y_{ij}$. In formulas, the algorithm we propose cycles over the following adaptive gradient steps:
@@
$\hat{\mathbf{G}}_{\textsc{u},{\scriptscriptstyle{I}}:}^{t}=(m/m_{{\scriptscriptstyle{J}}}^{\star})\,\accentset{\mbox{\large.}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t}\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}+\lambda\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t},$ $\hat{\mathbf{H}}_{\textsc{u},{\scriptscriptstyle{I}}:}^{t}=(m/m_{{\scriptscriptstyle{J}}}^{\star})\,\accentset{\mbox{\large..}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t}(\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}*\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t})+\boldsymbol{\Lambda},$ (15)
$\hat{\mathbf{G}}_{\textsc{v},{\scriptscriptstyle{J}}:}^{t}=(n/n_{{\scriptscriptstyle{I}}}^{\star})\,\accentset{\mbox{\large.}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t\top}\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}+\lambda\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t},$ $\hat{\mathbf{H}}_{\textsc{v},{\scriptscriptstyle{J}}:}^{t}=(n/n_{{\scriptscriptstyle{I}}}^{\star})\,\accentset{\mbox{\large..}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t\top}(\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}*\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t})+\boldsymbol{\Lambda}.$
Here, $(m/m_{{\scriptscriptstyle{J}}}^{\star})\,(\accentset{\mbox{\large.}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t}\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t})_{ih}=(m/m_{{\scriptscriptstyle{J}}}^{\star})\,\sum_{j\in J}\accentset{\mbox{\large.}}{D}_{ij}^{t}v_{jh}^{t}$ is an unbiased stochastic estimate of $(\accentset{\mbox{\large.}}{\mathbf{D}}^{t}\mathbf{V})_{ih}=\sum_{j=1}^{m}\accentset{\mbox{\large.}}{D}_{ij}^{t}v_{jh}^{t}$ for $i\in I$. Similarly, it is easy to show that the other minibatch averages in (15) are unbiased estimates of the corresponding batch quantities. Algorithm 2 provides a pseudo-code description of the proposed procedure, while Figure 1 provides a graphical representation of the updates.
If the dispersion parameter $\phi$ is unknown and has to be estimated from the data, we can also adopt a smoothed stochastic estimator obtained through exponential averaging. More details are provided in the Appendix.
@@
            Compute the subsampled likelihood derivatives
@@
             $\accentset{\mbox{\large..}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t}\leftarrow\mathbf{W}_{\scriptscriptstyle{B}}/\big{\{}\phi^{t}\,\nu(\boldsymbol{\mu}_{\scriptscriptstyle{B}}^{t})*\accentset{\mbox{\large.}}{g}(\boldsymbol{\mu}_{\scriptscriptstyle{B}}^{t})^{2}\big{\}}$; $O(n_{{\scriptscriptstyle{I}}}^{\star}m_{{\scriptscriptstyle{J}}}^{\star})$
            Compute the smoothed gradients and update $\mathbf{V}$
             $\hat{\mathbf{G}}_{\textsc{v},{\scriptscriptstyle{J}}:}^{t}\leftarrow(n/n_{\scriptscriptstyle{I}}^{*})\accentset{\mbox{\large.}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t\top}\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}-\lambda\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}$;  $\hat{\mathbf{H}}_{\textsc{v},{\scriptscriptstyle{J}}:}^{t}\leftarrow(n/n_{\scriptscriptstyle{I}}^{*})\accentset{\mbox{\large..}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t\top}(\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}*\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t})-\boldsymbol{\Lambda}$; $O(n_{{\scriptscriptstyle{I}}}^{\star}m_{{\scriptscriptstyle{J}}}^{\star}d)$
@@
            Compute the smoothed gradients and update $\mathbf{U}$
             $\hat{\mathbf{G}}_{\textsc{u},{\scriptscriptstyle{I}}:}^{t}\leftarrow(m/m_{\scriptscriptstyle{J}}^{*})\accentset{\mbox{\large.}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t}\,\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}-\lambda\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}$;  $\hat{\mathbf{H}}_{\textsc{u},{\scriptscriptstyle{I}}:}^{t}\leftarrow(m/m_{\scriptscriptstyle{J}}^{*})\accentset{\mbox{\large..}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t}\,(\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}*\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t})-\boldsymbol{\Lambda}$; $O(n_{{\scriptscriptstyle{I}}}^{\star}m_{{\scriptscriptstyle{J}}}^{\star}d)$

### v2

@@ 3.2.1 Improving naïve stochastic gradient
where $\bar{\mathbf{G}}_{\psi}^{t}$ and $\bar{\mathbf{H}}_{\psi}^{t}$ are smoothed estimates of the derivatives of (6), while $\alpha_{t}$ is a scalar bias-correction factor. Similar to Adam (Kingma and Ba,, 2014) and AMSGrad (Reddi et al.,, 2019), we update the smoothed gradients, $\bar{\mathbf{G}}_{\psi}^{t}$ and $\bar{\mathbf{H}}_{\psi}^{t}$, using an exponential moving average of the previous gradient values and the current stochastic estimates $\hat{\mathbf{G}}_{\psi}^{t}$ and $\hat{\mathbf{H}}_{\psi}^{t}$, namely
@@
with $\alpha_{1},\alpha_{2}\in(0,1]$ being user-specified smoothing coefficients, while $\alpha_{t}=(1-\alpha_{2}^{t})/(1-\alpha_{1}^{t})$ is introduced to filter out the bias induced by the exponential moving average in (12). Typical values for the smoothing coefficients are $\alpha_{1}=0.1$ and $\alpha_{2}=0.01$, respectively. Such a smoothing technique has two main advantages: it speeds up the convergence using the inertia accumulated by previous gradients and, at the same time, it stabilizes the optimization, reducing the noise around the gradient estimate. We refer the reader to Kingma and Ba, (2014) and Reddi et al., (2019) for a deeper discussion on the benefits of bias corrected exponential gradient averaging in high-dimensional stochastic optimization problems.
Different from the original implementation of Adam (Kingma and Ba,, 2014), which only involves first-order information, we scale the smoothed gradient estimate using a smoothed diagonal Hessian approximation. This strategy allows us to directly extend the diagonal quasi-Newton algorithm of Kidziński et al., (2022) in a stochastic vein without increasing the computational overload or the stability of the optimization. Indeed, the derivatives of the deviance function of a GLM model are always well-defined, bounded away from zero, available in closed form and easy to compute. This is not the case for deep neural networks and other machine learning models for which Adam and AMSGrad have been originally developed.
Analogously to the quasi-Newton update (10) for GMF models, also (11) can be written in the block-wise form
@@
In this formulation, update (10) and (13) have a computational complexity proportional to the dimension of the matrices $\mathbf{U}$ and $\mathbf{V}$, namely $O((n+m)d)$. Since the dimension of the parameter space can not be further compressed without losing prediction power, the alternative way to speed up the computations is to efficiently approximate the gradients, $\hat{\mathbf{G}}_{\psi}^{t}=(\hat{\mathbf{G}}_{\textsc{u}}^{t},\hat{\mathbf{G}}_{\textsc{v}}^{t})$ and $\hat{\mathbf{H}}_{\psi}^{t}=(\hat{\mathbf{H}}_{\textsc{u}}^{t},\hat{\mathbf{H}}_{\textsc{v}}^{t})$.
@@
In what follows, we discuss a new stochastic optimization method for GMF models based on formula (13), which uses local parameter updates along with minibatch estimates of the current gradients to reduce the computational complexity of the resulting algorithm. To this end, we introduce the following notation: $I=\{i_{1},\dots,i_{n_{{\scriptscriptstyle{I}}}^{\star}}\}\subseteq\{1,\dots,n\}$ is a subset of row-indices of dimension $n_{{\scriptscriptstyle{I}}}^{\star}$, $J=\{j_{1},\dots,j_{m_{{\scriptscriptstyle{J}}}^{\star}}\}\subseteq\{1,\dots,m\}$ is a subset of column-indices of dimension $m_{{\scriptscriptstyle{J}}}^{\star}$, $B=I\times J$ is the Cartesian product between $I$ and $J$. Finally, $\mathbf{Y}_{{\scriptscriptstyle{I}}:}=\{\mathbf{y}_{i:}\}_{i\in I}$, $\mathbf{Y}_{:{\scriptscriptstyle{J}}}=\{\mathbf{y}_{:j}\}_{j\in J}$ and $\mathbf{Y}_{{\scriptscriptstyle{B}}}=\{y_{ij}\}_{(i,j)\in B}$ denote the corresponding sub-matrices, also called row-, column- and block-minibatch subsamples of the original data matrix, respectively.
@@ 3.2.2 Block-wise adaptive stochastic gradient descent
Both the exact and stochastic quasi-Newton methods identified by equations (10) and (13) entirely update $\mathbf{U}$ and $\mathbf{V}$ at each iteration. Despite being an effective and parallelizable strategy, in many applicative contexts, when it comes to factorizing huge matrices, an entire update of the parameters could be extremely expensive in terms of memory allocation and execution time. This is a well-understood issue in the literature on recommendation systems, where standard matrix completion problems may involve matrices with millions of rows and columns; see, e.g., Koren et al., (2009), Mairal et al., (2010), Recht and Ré, (2013), Mensch et al., (2017). Moreover, batch optimization strategies do not generalize well to stream data contexts, where the data arrive sequentially and the parameters must be updated on-the-fly as new sets of observations come in.
A classic solution is to perform iterative element-wise SGD steps using only one entry of the data matrix, $y_{ij}$, at each iteration, thus updating the low-rank decomposition matrices row-by-row through the paired equations $\mathbf{u}_{i:}^{t+1}\leftarrow\mathbf{u}_{i:}^{t}-\rho_{t}\mathbf{G}_{\textsc{u},i:}^{t}$ and $\mathbf{v}_{j:}^{t+1}\leftarrow\mathbf{v}_{j:}^{t}-\rho_{t}\mathbf{G}_{\textsc{v},j:}^{t}$. In the matrix factorization literature, this strategy is also known as online SGD algorithm, since it permits updating the parameter estimates dynamically when a new observation is gathered.
To stabilize the optimization and speed up the convergence, we generalize the online SGD approach in two directions: we consider local stochastic quasi-Newton updates in place of naïve gradient steps, and we use block-wise minibatches of the original data matrix, $\mathbf{Y}_{\scriptscriptstyle{B}}$, instead of singletons, $y_{ij}$. In formulas, the algorithm we propose cycles over the following adaptive gradient steps:
@@
$\hat{\mathbf{G}}_{\textsc{u},{\scriptscriptstyle{I}}:}^{t}=(m/m_{{\scriptscriptstyle{J}}}^{\star})\,\accentset{\mbox{\large.}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t}\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}+\lambda\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t},$ $\hat{\mathbf{H}}_{\textsc{u},{\scriptscriptstyle{I}}:}^{t}=(m/m_{{\scriptscriptstyle{J}}}^{\star})\,\accentset{\mbox{\large.\hskip-1.33333pt.}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t}(\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}*\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t})+\boldsymbol{\Lambda},$ (15)
$\hat{\mathbf{G}}_{\textsc{v},{\scriptscriptstyle{J}}:}^{t}=(n/n_{{\scriptscriptstyle{I}}}^{\star})\,\accentset{\mbox{\large.}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t\top}\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}+\lambda\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t},$ $\hat{\mathbf{H}}_{\textsc{v},{\scriptscriptstyle{J}}:}^{t}=(n/n_{{\scriptscriptstyle{I}}}^{\star})\,\accentset{\mbox{\large.\hskip-1.33333pt.}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t\top}(\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}*\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t})+\boldsymbol{\Lambda}.$
Here, $n_{{\scriptscriptstyle{I}}}^{\star}$ and $m_{{\scriptscriptstyle{J}}}^{\star}$ denote the number of rows and columns of each minibatch and $(m/m_{{\scriptscriptstyle{J}}}^{\star})\,(\accentset{\mbox{\large.}}{\mathbf{D}}_{{\scriptscriptstyle{B}}}^{t}\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t})_{ih}=(m/m_{{\scriptscriptstyle{J}}}^{\star})\,\sum_{j\in J}\accentset{\mbox{\large.}}{D}_{ij}^{t}v_{jh}^{t}$ is an unbiased stochastic estimate of $(\accentset{\mbox{\large.}}{\mathbf{D}}^{t}\mathbf{V})_{ih}=\sum_{j=1}^{m}\accentset{\mbox{\large.}}{D}_{ij}^{t}v_{jh}^{t}$ for $i\in I$. Similarly, it is easy to show that the other minibatch averages in (15) are unbiased estimates of the corresponding batch quantities. Figure 1 provides a graphical representation of the updates, while Algorithm 1 provides a pseudo-code description of the proposed procedure. Overall, each iteration of Algorithm 1 requires $O((n_{{\scriptscriptstyle{I}}}^{\star}+m_{{\scriptscriptstyle{J}}}^{\star})d+n_{{\scriptscriptstyle{I}}}^{\star}m_{{\scriptscriptstyle{J}}}^{\star}d)$ floating point operations and $O((n_{{\scriptscriptstyle{I}}}^{\star}+m_{{\scriptscriptstyle{J}}}^{\star})d+n_{{\scriptscriptstyle{I}}}^{\star}m_{{\scriptscriptstyle{J}}}^{\star})$ memory allocations.
Figure 1: Graphical representation of the stochastic gradient updates employed at the $t$th iteration of Algorithm 1. Left: generalized matrix factorization model (3). Middle: updates of the penalized log-likelihood gradients (15). Right: adaptive gradient step (14). The colored and empty cells highlight the sub-sampled data used and not used at the $t$th update, respectively. To save space, the calculations of the second-order differentials and the gradient smoothing are not displayed here.
@@ 1. Sample the minibatch set
     2. Compute the subsampled likelihood derivatives
@@
       $\accentset{\mbox{\large.\hskip-1.33333pt.}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t}\leftarrow\mathbf{W}_{\scriptscriptstyle{B}}/\big\{\phi^{t}\,\nu(\boldsymbol{\mu}_{\scriptscriptstyle{B}}^{t})*\accentset{\mbox{\large.}}{g}(\boldsymbol{\mu}_{\scriptscriptstyle{B}}^{t})^{2}\big\}$; $O(n_{{\scriptscriptstyle{I}}}^{\star}m_{{\scriptscriptstyle{J}}}^{\star})$
     3. Compute the smoothed gradients and update $\mathbf{V}$
        $\hat{\mathbf{G}}_{\textsc{v},{\scriptscriptstyle{J}}:}^{t}\leftarrow(n/n_{\scriptscriptstyle{I}}^{*})\accentset{\mbox{\large.}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t\top}\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}-\lambda\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}$;  $\hat{\mathbf{H}}_{\textsc{v},{\scriptscriptstyle{J}}:}^{t}\leftarrow(n/n_{\scriptscriptstyle{I}}^{*})\accentset{\mbox{\large.\hskip-1.33333pt.}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t\top}(\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}*\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t})-\boldsymbol{\Lambda}$; $O(n_{{\scriptscriptstyle{I}}}^{\star}m_{{\scriptscriptstyle{J}}}^{\star}d)$
@@ 3. Compute the smoothed gradients and update $\mathbf{V}$
     4. Compute the smoothed gradients and update $\mathbf{U}$
        $\hat{\mathbf{G}}_{\textsc{u},{\scriptscriptstyle{I}}:}^{t}\leftarrow(m/m_{\scriptscriptstyle{J}}^{*})\accentset{\mbox{\large.}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t}\,\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}-\lambda\mathbf{U}_{{\scriptscriptstyle{I}}:}^{t}$;  $\hat{\mathbf{H}}_{\textsc{u},{\scriptscriptstyle{I}}:}^{t}\leftarrow(m/m_{\scriptscriptstyle{J}}^{*})\accentset{\mbox{\large.\hskip-1.33333pt.}}{\mathbf{D}}_{\scriptscriptstyle{B}}^{t}\,(\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t}*\mathbf{V}_{{\scriptscriptstyle{J}}:}^{t})-\boldsymbol{\Lambda}$; $O(n_{{\scriptscriptstyle{I}}}^{\star}m_{{\scriptscriptstyle{J}}}^{\star}d)$